Lledtech Corporation is expanding rapidly and currently needs to retain all its earnings; hence, it does not pay dividends. However, investors expect Lledtech to begin paying dividends, beginning with a dividend of $1.75 coming 3 years from today. The dividend should grow rapidly-at a rate of 30% per year-during Years 4 and 5; but after Year 5, growth should be a constant 5% per year. If the required return on Lledtech is 13%, what is the value of the stock today?

Answer: $25.38

Please show all work and formulas used for a financial calculator, not excel.
Answer #1

D3=1.75

D4=(1.75*1.3)=2.275

D5=(2.275*1.3)=2.9575

Value after year 5=(D5*Growth Rate)/(Required rate-Growth Rate)

=(2.9575*1.05)/(0.13-0.05)

=$38.8171875

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.75/1.13^3+2.275/1.13^4+2.9575/1.13^5+$38.8171875/1.13^5

which is equal to

=$25.38(Approx).